Calorimeter data quality survey

In this page are listed all the hardware problems identified during the analysis of calorimeter data (noise, LED or cosmics)


Problematic channels

List

The problematic channels listed below have been identified thanks to the analysis of Cosmics and LED data taken in february 2008.
Detailed analysis of the data can be found here and there.

Irina's survey summary (beg. march 2008) can be found here.

Calo Crate FEB RO channel diagnostic Data seen fixed
Hcal 22 10 30 dead LED/Cosmics 13/02/08  
Hcal 23 8 28 dead LED/Cosmics 13/02/08  
Hcal 22 13 [PIN] 2 dead LED 13/02/08  
Ecal 13 2 7 dead LED 13/03/08  
Ecal 8 5 29 dead LED/Cosmics 13/02/08 20/02/08 PM replaced
Ecal 10 0 21 dead LED/Cosmics 13/02/08 20/02/08 PM replaced
Ecal 11 7 12 dead LED/Cosmics 13/02/08 20/02/08 PM replaced
Ecal 12 2 1 dead LED/Cosmics 13/02/08 13/03/08 no longer pb(?)
Ecal 14 8 6 dead LED/Cosmics 13/02/08 13/03/08 no longer pb(?)
Ecal 9 8 31 dead (FEB pb to be fixed) LED/Cosmics 13/02/08  
Ecal 14 9 24 dead LED/Cosmics 13/02/08 20/02/08 bad connector
Ecal 12 2->11 3,11,19,27 low signal LED/Cosmics 13/02/08 20/02/08 power line pb
Ecal 13 2->4 7,15,23 low signal LED/Cosmics 13/02/08 20/02/08 power line pb
Ecal 14 11->13 7,15,23 low signal LED/Cosmics 13/02/08 20/02/08 power line pb
Ecal 11 8 4->7, 12->15, 20->23, 28->31 low signal (bad timing ?) LED 13/02/08  
Ecal 11 9 0->3, 8->11, 16->19, 24->27 low signal (bad timing ?) LED 13/02/08  
Ecal 11 7 12 hot Cosmics 02/2008 20/02/08 found dead - PMT replaced
Ecal 13 12 9 hot Cosmics 02/2008 20/02/08 found dead - PMT replaced
Ecal 14 7 28 very hot Cosmics/High trigger rate 13/02/08 20/02/08 found dead - PMT replaced
Ecal 14 8 7 & 15 & 23 & 31 very hot Cosmics/High trigger rate 13/02/08  
Ecal 14 9 13 hot Cosmics/High trigger rate 13/02/08  
Ecal 14 9 21 hot Cosmics/High trigger rate 13/02/08 20/02/08 found dead - PMT replaced
Hcal 22 8 6-7 & 14-15 & 22-23 & 30-31 hot Cosmics 13/02/08  
Hcal 23 6 31 hot Cosmics 13/02/08  
Hcal 23 7 10 & 18 & 26 hot Cosmics 13/02/08  

Figures

  • DEAD AND LOW SIGNAL CHANNELS : Ecal (right) LED scan : ADC (summed over 5 consecutive BX) as a function of FEB and channel
    hcal_quality.jpg

  • DEAD CHANNELS : Hcal (left) and Ecal (right) LED scan : ADC (summed over 3 consecutive BX) as a function of FEB and channel
    hcal_quality.jpg ecal_quality.jpg

  • HOT CHANNELS : Hcal (left) and Ecal (right) L0 trigger : ADC as a function of FEB and channel
    hcal_hot_L0.jpg ecal_hot_L0.jpg

  • HOT Ecal channels
    hot_ecal_channels.jpg


DAQ problems

List

  • problem of transmission with ecTell07 (time to time wrong header and/or wrong L0ADC/ADC)
  • 1 uncorrect L0ADC/ADC observed with ecTell04
  • problem with FEB#6 in Prs0 (psTell01) : wrong header with CrateID/FEBId=0
  • problem with 4 FEB in Prs2 (FEB#4,5,6,7) --> faulty PGA of psTell03 ?

Figures :

  • PRS C-side
    prs0123_22Feb08_T0_Prev1-zoom.JPG
  • L0ADC/ADC matching
    ecal_l0adc.jpg


Wrong settings

List

  • Ecal and Hcal PIN-diode signal is shifted ~ -2 BX with respect to PMT signal
  • Ecal PIN-diode signal for BOX#5 and BOX#6 additional ~ -2BX delay (10m cable vs 17m cable)
  • CROC & TELL1 BCIDs not aligned -> production of error banks : BCID offset to be set

-- OlivierDeschamps - 26 Feb 2008

Topic attachments
I Attachment History Action Size Date Who Comment
Microsoft Word filedoc ECALbadChan.doc r1 manage 411.0 K 2008-03-14 - 10:11 OlivierDeschamps  
JPEGjpg ecal_hot_L0.jpg r1 manage 55.1 K 2008-02-27 - 12:43 OlivierDeschamps  
JPEGjpg ecal_l0adc.jpg r1 manage 26.9 K 2008-02-27 - 12:46 OlivierDeschamps  
JPEGjpg ecal_quality.jpg r1 manage 154.0 K 2008-02-27 - 12:41 OlivierDeschamps  
JPEGjpg ecal_quality_031308.jpg r1 manage 147.0 K 2008-03-13 - 13:22 OlivierDeschamps  
JPEGjpg hcal_hot_L0.jpg r1 manage 50.8 K 2008-02-27 - 12:42 OlivierDeschamps  
JPEGjpg hcal_quality.jpg r1 manage 150.4 K 2008-02-27 - 12:41 OlivierDeschamps  
JPEGjpg hot_ecal_channels.jpg r2 r1 manage 23.8 K 2008-02-27 - 13:05 OlivierDeschamps  
JPEGjpg prs0123_22Feb08_T0_Prev1-zoom.JPG r1 manage 36.5 K 2008-02-27 - 12:45 OlivierDeschamps  
Edit | Attach | Watch | Print version | History: r8 < r7 < r6 < r5 < r4 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r8 - 2008-06-19 - OlivierDeschamps
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LHCb All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback